# Nested subspace learning with flags

### v1

@@ 4.2.1 Application of the flag trick to RSR
Among the large family of methods for robust subspace recovery (Lerman and Maunu, 2018b) we consider the one of least absolute deviation (LAD) minimization, which can be explicitly formulated as an optimization problem on Grassmannians. PCA minimizes the sum of squared Euclidean distances between the points and the subspace. In the case of an outlier-contaminated dataset, the squared Euclidean distances might penalize too much the outliers, and therefore make the optimal subspace too much influenced by the outliers. To circumvent this well-known sensitivity of squared norms to outliers, many works propose to minimize the sum of absolute Euclidean distances, which defines the LAD minimization problem:
$\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,\sum_{i=1}^{n}\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}.$ (12)
The latter has the interest of being rotationally invariant (Ding et al., 2006) but the drawback of being NP-hard (McCoy and Tropp, 2011; Lerman and Maunu, 2018b) and obviously non-convex since Grassmannians are not. A first body of works relaxes the problem, for instance by optimizing on the convex hull of Grassmannians (McCoy and Tropp, 2011; Xu et al., 2012; Zhang and Lerman, 2014; Lerman et al., 2015). A second body of works directly optimizes the LAD on Grassmannians, either with an IRLS algorithm (Lerman and Maunu, 2018a) or with a geodesic gradient descent (Maunu et al., 2019), both achieving very good results in terms of recovery and speed. The following proposition applies the flag trick to the LAD problem (12).
@@
The flag trick applied to LAD reads
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\sum_{i=1}^{n}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2},$ (13)
and is equivalent to the following optimization problem:
$\underset{U_{1:d+1}\in\mathcal{O}(p)}{\operatorname{argmin}}\,\sum_{i=1}^{n}\sqrt{\sum_{k=1}^{d+1}\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}}.$ (14)
Proof  The proof is given in Appendix (Appendix A).
Equation (14) tells us several things. First of all, due to the $((k-1)/d)^{2}$ weights, the points that are far from the center should be in the first principal subspaces. Second, the square root decreases the influence of the outlying points, compared to the nested PCA of Theorem 4. Third, although less obvious, we can see that numerical issues are less prone to happen with the flag trick (14) than with the LAD (12), since the quantity under the square root is zero only when the first subspace $\leavevmode\nobreak\ {\mathcal{S}_{1}=\operatorname{Span}(U_{1})}$ contains a data point. Therefore, whenever $q_{1}$ is smaller than the dimension $q$ that one would have tried for classical RSR, the non-differentiability and exploding-gradient issues caused by the square root are less likely. Finally, since the nested LAD minimization (13) is nothing but a robust version of the nested PCA of Theorem 4 ($\sum_{i=1}^{n}\|x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\|_{2}^{2}$), a natural idea can be to initialize the optimization algorithm with the nested PCA solution. This is what is done in Maunu et al. (2019) for LAD minimization, and it is coming with stronger recovery guarantees.
@@ 4.2.2 Nestedness experiments for RSR
[…]
Figure 2: Illustration of the nestedness issue in robust subspace recovery. Given a dataset consisting in a mixture of inliers (blue) and outliers (orange) we plot its projection onto the optimal 1D subspace and 2D subspace obtained by solving the associated Grassmannian optimization problem (12) or flag optimization problem (13). We can see that the Grassmann representations are not nested, while the flag representations are nested and robust to outliers.
Figure 3: Distributions of (increasing) Euclidean reconstruction errors $\|x_{i}-\Pi_{\mathcal{S}_{1:d}^{*}}x_{i}\|_{i=1\dots n}$ on the corrupted digits dataset for Grassmann and flag methods. While inliers and outliers are mixed with the subspace method, we can see a clear transition with flags. This can be explained by the multilevel nature of the flag trick.
@@
Second, we perform an outlier detection experiment. A common methodology to detect outliers in a corrupted dataset is to first look for an outlier-robust subspace and then plot the distribution of distances between the data points and their projection onto the recovered subspace. This distribution is expected to show a clear gap between the inliers and outliers (Vidal et al., 2016, Fig. 3.7). However, in practice, one does not know which subspace dimension $q$ to choose. If $q$ is too large, then the recovered subspace may contain both inliers and outliers, and therefore the distribution of distances might be roughly 00. In contrast, if $q$ is too small, then some inliers may lie too far from the recovered subspace and be detected as outliers. An idea in the spirit of the flag trick is to perform an average ensembling of the reconstruction errors. More specifically, if $\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}$ is the classical measure of robust reconstruction error, then we compute $\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$. Such a score extracts information from projections at different levels and might result in a richer multilevel analysis. We consider a dataset where the inliers are images of $8\times 8$ handwritten 00’s and outliers correspond to other digits from $1$ to $9$, all extracted from the classical handwritten digits dataset (Alpaydin and Kaynak, 1998). The ambient dimension is $p=64$, the number of inliers is $n_{\mathrm{in}}=90$ and the number of outliers is $n_{\mathrm{out}}=10$. The intrinsic dimensions that we try are $q_{1:3}=(1,2,5)$. We plot the reconstruction error for the points of the digits dataset on the optimal flag $\mathcal{S}_{1:3}^{*}\in\operatorname{Fl}(p,(1,2,5))$ in Figure 3. We compare it to the metric on $\mathcal{S}_{3}\in\operatorname{Gr}(p,q_{3})$. We can see that the flag trick enables to clearly distinguish the inliers from the outliers compared to the Grassmann-based method, which is a consequence of the multilevel nature of flags.
@@ 4.2.3 Discussion on RSR optimization and objective functions
In all the experiments of this paper, we use a steepest descent method on flag manifolds (Algorithm 1) to solve the flag problems. However, for the specific problem of RSR (13), we believe that more adapted algorithms should be derived, notably due to the non-differentiability and exploding-gradient issues caused by the square root. To that extent, we derive in appendix (Appendix A) an IRLS scheme (Algorithm 3) for RSR. In short, the RSR problem (13) can be reformulated as a weighted least squares problem $\sum_{i=1}^{n}w_{i}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}^{2}$ with $w_{i}={1}/{\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}}$ and optimized iteratively, with explicit expressions obtained via our central Theorem 4. We insist on the fact that such an IRLS algorithm could not be developed with the flagification of Pennec (2018); Mankovich et al. (2024), since a sum of square roots does not correspond to a least-squares problem.
@@
In this work, we explore one specific problem of RSR for conciseness, but we could investigate many other related problems, including robust PCA. Notably, drawing from the Grassmann averages (GA) method (Hauberg et al., 2016), one could develop many new multilevel RSR and RPCA objective functions. The idea behind GA is to replace data points with 1D subspaces ($\mathcal{S}_{i}=\operatorname{Span}(x_{i})$) and then perform subspace averaging methods to find a robust prototype for the dataset. GA ends up solving problems of the form $\operatorname{argmin}_{\mathcal{S}\in\operatorname{Gr}(p,1)}\sum_{i=1}^{n}w_{i}\,\operatorname{dist}_{\operatorname{Gr}(p,1)}^{2}(\operatorname{Span}(x_{i}),\mathcal{S})$, where $w_{i}$ are some weights and $\operatorname{dist}_{\operatorname{Gr}(p,1)}$ is a particular subspace distance detailed in Hauberg et al. (2016). Using instead some multidimensional subspace distances, like the principal angles and its variants (Hamm and Lee, 2008; Ye and Lim, 2016), we can develop many variants of the Grassmann averages, of the form $\operatorname{argmin}_{\mathcal{S}_{1:d}\in\operatorname{Gr}(p,q)}\sum_{i=1}^{n}w_{i}\,\rho(\sqrt{{x_{i}}^{\top}\Pi_{\mathcal{S}}{x_{i}}})$, where $\rho\colon\mathbb{R}\to\mathbb{R}$ is a real function, like $\rho(x)=\arccos(x)$ if we want subspace-angle-like distances, $\rho(x)=-x^{2}$ if we want PCA-like solutions, and many other possible robust variants. Applying the flag trick to those problems yields the following robust multilevel problem: $\operatorname{argmin}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\sum_{i=1}^{n}w_{i}\rho(\sqrt{{x_{i}}^{\top}\Pi_{\mathcal{S}_{1:d}}{x_{i}}})$.
@@ 4.3 The flag trick for trace ratio problems
Trace ratio problems are ubiquitous in machine learning (Ngo et al., 2012). They write as:
$\underset{U\in\operatorname{St}(p,q)}{\operatorname{argmax}}\,\frac{\operatorname{tr}\left(U^{\top}AU\right)}{\operatorname{tr}\left(U^{\top}BU\right)},$ (15)
where $A,B\in\mathbb{R}^{p\times p}$ are positive semi-definite matrices, with $\operatorname{rank}(B)>p-q$.
@@
A famous example of TR problem is Fisher’s linear discriminant analysis (LDA) (Fisher, 1936; Belhumeur et al., 1997). It is common in machine learning to project the data onto a low-dimensional subspace before fitting a classifier, in order to circumvent the curse of dimensionality. It is well known that performing an unsupervised dimension reduction method like PCA comes with the risks of mixing up the classes, since the directions of maximal variance are not necessarily the most discriminating ones (Chang, 1983). The goal of LDA is to use the knowledge of the data labels to learn a linear subspace that does not mix the classes. Let $\leavevmode\nobreak\ {X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}}$ be a dataset with labels $Y:=[y_{1}|\dots|y_{n}]\in{[1,C]}^{n}$. Let $\mu=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ be the dataset mean and $\mu_{c}=\frac{1}{\#\{i:y_{i}=c\}}\sum_{i:y_{i}=c}x_{i}$ be the class-wise means. The idea of LDA is to search for a subspace $\mathcal{S}\in\operatorname{Gr}(p,q)$ that simultaneously maximizes the projected between-class variance $\sum_{c=1}^{C}\|\Pi_{\mathcal{S}}\mu_{c}-\Pi_{\mathcal{S}}\mu\|_{2}^{2}$ and minimizes the projected within-class variance $\sum_{c=1}^{C}\sum_{i:y_{i}=c}\|\Pi_{\mathcal{S}}x_{i}-\Pi_{\mathcal{S}}\mu_{c}\|_{2}^{2}$. This can be reformulated as a trace ratio problem (15), with $A=\sum_{c=1}^{C}(\mu_{c}-\mu)(\mu_{c}-\mu)^{\top}$ and $B=\sum_{c=1}^{C}\sum_{i:y_{i}=c}(x_{i}-\mu_{c})(x_{i}-\mu_{c})^{\top}$.
@@
More generally, a large family of dimension reduction methods can be reformulated as a TR problem. The seminal work of Yan et al. (2007) shows that many dimension reduction and manifold learning objective functions can be written as a trace ratio involving Laplacian matrices of attraction and repulsion graphs. Intuitively, those graphs determine which points should be close in the latent space and which ones should be far apart. Other methods involving a ratio of traces are multi-view learning (Wang et al., 2023), partial least squares (PLS) (Geladi and Kowalski, 1986; Barker and Rayens, 2003) and canonical correlation analysis (CCA) (Hardoon et al., 2004), although these methods are originally sequential problems (cf. footnote 1) and not subspace problems.
@@
Classical Newton-like algorithms for solving the TR problem (15) come from the seminal works of Guo et al. (2003); Wang et al. (2007); Jia et al. (2009). The interest of optimizing a trace-ratio instead of a ratio-trace (of the form $\operatorname{tr}\left((U^{\top}BU)^{-1}(U^{\top}AU)\right)$), that enjoys an explicit solution given by a generalized eigenvalue decomposition, is also tackled in those papers. The repulsion Laplaceans (Kokiopoulou and Saad, 2009) instead propose to solve a regularized version $\operatorname{tr}\left(U^{\top}BU\right)-\rho\operatorname{tr}\left(U^{\top}AU\right)$, which enjoys a closed-form, but has a hyperparameter $\rho$, which is directly optimized in the classical Newton-like algorithms for trace ratio problems.

### v2

@@ 4.2.1 Application of the Flag Trick to RSR
Among the large family of methods for robust subspace recovery presented in lerman_overview_2018, we consider the one of least absolute deviation (LAD):
$\operatorname*{arg\,min}_{\mathcal{S}\in\operatorname{Gr}(p,q)}\sum_{i=1}^{n}\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}.$ (12)
It is motivated by the sensitivity of squared norms to outliers, which makes PCA (2) unsuitable for outlier-contaminated data. Problem (12) has the advantage of being rotationally invariant (ding_r1-pca_2006) but the drawback of being NP-hard (mccoy_two_2011; lerman_overview_2018) and obviously non-convex—since Grassmannians are not. A first body of works relaxes the problem, for instance by optimizing on the convex hull of Grassmannians (mccoy_two_2011; xu_robust_2012; zhang_novel_2014; lerman_robust_2015). A second body of works directly optimizes the LAD criterion on Grassmannians, either with an IRLS algorithm (lerman_fast_2018) or with a geodesic gradient descent (maunu_well-tempered_2019), both achieving very good results in terms of recovery and speed. The following proposition applies the flag trick to the LAD problem (12).
@@
The flag trick transforms the least absolute deviation problem (12) into:
$\operatorname*{arg\,min}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\sum_{i=1}^{n}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2},$ (13)
which is equivalent to the following optimization problem:
$\operatorname*{arg\,min}_{U_{1:d}\in\operatorname{St}(p,q)}\sum_{i=1}^{n}\sqrt{\left\lVert x_{i}\right\rVert_{2}^{2}-\left\lVert{U_{1:d}}^{\top}x_{i}\right\rVert_{2}^{2}+\sum_{k=1}^{d}\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}}.$ (14)
Proof The proof is given in Appendix A.
Remark 15 (Generalization of RSR)
We can check that Equation (14) generalizes the original RSR problem (12) from maunu_well-tempered_2019. Indeed, if we consider the particular case $d=1$, $U\coloneqq U_{1:d}$, $\mathcal{S}\coloneqq\operatorname{Span}(U_{1})$ (Grassmannian), then we get:
$\sum_{i=1}^{n}\sqrt{\left\lVert x_{i}\right\rVert_{2}^{2}-\left\lVert{U_{1:d}}^{\top}x_{i}\right\rVert_{2}^{2}+\sum_{k=1}^{d}\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}}=\sum_{i=1}^{n}\sqrt{\left\lVert x_{i}\right\rVert_{2}^{2}-\left\lVert{U_{1:d}}^{\top}x_{i}\right\rVert_{2}^{2}+0}=\\ \sum_{i=1}^{n}\left\lVert(I_{p}-UU^{\top})x_{i}\right\rVert_{2}=\sum_{i=1}^{n}\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}.$ (15)
Hence, (14) is a multilevel generalization of the RSR problem from maunu_well-tempered_2019. The new term under the square root, $\sum_{k=1}^{d}\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}$, is nonzero when $d>1$ (i.e., for flag manifolds refining Grassmannians). This additional term corresponds to a weighted sum over the blocks of the orthonormal frame $U_{1:d}\in\operatorname{St}(p,q)$, with increasing weights $((k-1)/d)^{2}$. Therefore, intuitively, the variance in $\mathrm{Im}(U_{k}^{*})$ is expected to decrease with $k$. In other words, the points that are far from the center—i.e., with large norm—should be in the first principal subspaces. Hence, we can interpret our new criterion as a robust generalization of the nested PCA (Theorem 8), with a least influence of outliers due to the square root.
Moreover, we can see that numerical issues are less prone to occur with the flag-tricked problem (13) than with the original one (12). Indeed, the numerical issues of the square root arise when the radicand ${\|x_{i}-\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}x_{i}\|}_{2}^{2}$ is close to zero. By nestedness, this is the case if and only if $\mathcal{S}_{1}$ is close to $x_{i}$. Consequently, while numerical issues arise in classical RSR (12) whenever the subspace $\mathcal{S}$ is close to any data point, they arise in flag-tricked RSR (13) whenever both the subspace $\mathcal{S}$ and the smaller nested subspaces are close to that data point. Hence, whenever $q_{1}\coloneqq\min q_{1:d}$ is smaller than the dimension $q$ that one would have tried for classical RSR, the non-differentiability and exploding-gradient issues caused by the square root are less likely to occur.
Finally, since the flag-tricked problem (13) is nothing but a robust version of the nested PCA of Theorem 8 ($\sum_{i=1}^{n}{\|x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\|}_{2}^{2}$), a natural idea can be to initialize the optimization algorithm with the nested PCA solution. This is what is done in maunu_well-tempered_2019 for the original problem (12), and it is coming with exact recovery guarantees.
@@ 4.2.2 Nestedness Experiments for RSR
We first consider a dataset consisting in a mixture of two multivariate Gaussians: the inliers, with zero mean, covariance matrix $\operatorname{diag}\left(5,1,.1\right)$, $n_{\mathrm{in}}=450$ and the outliers, with zero mean, covariance matrix $\operatorname{diag}\left(.1,.1,5\right)$, $n_{\mathrm{out}}=50$. The dataset is therefore following the generalized haystack model of maunu_well-tempered_2019. The ambient dimension is $p=3$ and the intrinsic dimensions that we try are $q_{1:2}=(1,2)$. We run Algorithm 1 on Grassmann manifolds to solve the LAD problem (12), successively for $q_{1}=1$ and $q_{2}=2$. Then we plot the projections of the data points onto the optimal subspaces. We compare them to the nested projections onto the optimal flag output by running Algorithm 1 on $\operatorname{Fl}(3,(1,2))$ to solve (13). The results are shown in Figure 3.
Figure 3: Illustration of the nestedness issue in robust subspace recovery. Given a dataset consisting in a mixture of inliers (blue) and outliers (red) we plot its projection onto the optimal 1D subspace and 2D subspace obtained by solving the associated Grassmannian optimization problem (12) or flag optimization problem (13). We can see that the Grassmann representations are not nested, while the flag representations are nested and robust to outliers.
Figure 4: Euclidean reconstruction errors (sorted in ascending order) on the corrupted digits dataset for robust subspace recovery ($\|x_{i}-\Pi_{\mathcal{S}^{*}}x_{i}\|$, left) and its flag-tricked version ($\|x_{i}-\Pi_{\mathcal{S}_{1:d}^{*}}x_{i}\|$, right). With the Grassmann-based method, the distributions of reconstruction errors for the inliers and outliers intersect, meaning that we cannot fully distinguish the inliers from the outliers (see the gray zone on the left plot). In contrast, with the flag-based method, the distributions of reconstruction errors for the inliers and outliers are clearly separated, meaning that we can easily distinguish the inliers from the outliers (see the white zone on the right plot). This phenomenon can be explained by the multilevel nature of the flag trick (see the main text for more details).
@@
Second, we perform an outlier detection experiment. A common methodology to detect outliers in a corrupted dataset is to first look for an outlier-robust subspace and then plot the distribution of distances between the data points and their projection onto the recovered subspace. This distribution is expected to show a clear gap between the inliers and outliers (vidal_generalized_2016, Fig. 3.7). However, in practice, one does not know which subspace dimension $q$ to choose. If $q$ is too large, then the recovered subspace may contain both inliers and outliers, and therefore the distribution of distances might be roughly 0. In contrast, if $q$ is too small, then some inliers may lie too far from the recovered subspace and be detected as outliers. An idea in the spirit of the flag trick is to perform an average ensembling of the reconstruction errors. More specifically, if $\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}$ is the classical measure of robust reconstruction error, then we compute $\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$. Such a score extracts information from projections at different levels and might result in a richer multilevel analysis. We consider a dataset where the inliers are images of $8\times 8$ handwritten 0’s and outliers correspond to other digits from $1$ to $9$, all extracted from the classical handwritten digits dataset (alpaydin_optical_1998). The ambient dimension is $p=64$, the number of inliers is $n_{\mathrm{in}}=90$ and the number of outliers is $n_{\mathrm{out}}=10$. The intrinsic dimensions that we try are $q_{1:3}=(1,2,5)$. We plot the distribution of reconstruction errors for the points of the digits dataset on the optimal flag $\mathcal{S}_{1:3}^{*}\in\operatorname{Fl}(p,(1,2,5))$ in Figure 4. We compare it to the distribution of reconstruction errors on $\mathcal{S}^{*}\in\operatorname{Gr}(p,5)$. We can see that the flag trick enables to clearly distinguish the inliers from the outliers compared to the Grassmann-based method, which is a consequence of the multilevel nature of flags.
@@ 4.2.3 Discussion on RSR Optimization and Objective Functions
To substantiate the generality of the flag trick, we propose two extensions of the previously proposed multilevel RSR method: an IRLS optimization algorithm and a collection of new multilevel criteria. The experimentation of these extensions is left to future work since the goal is rather to show how the flag trick principle can be easily applied in different contexts.
@@
In all the experiments of this paper, we use a steepest descent method on flag manifolds (Algorithm 1) to solve the flag problems. However, for the specific problem of RSR (13), we believe that more adapted algorithms should be derived, notably due to the non-differentiability and exploding-gradient issues caused by the square root. To that extent, we derive in Appendix A an IRLS scheme (Algorithm 3) for RSR. In short, the RSR problem (13) can be reformulated as a weighted least squares problem $\sum_{i=1}^{n}w_{i}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}^{2}$ with $w_{i}={1}/{\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}}$ and optimized iteratively, with explicit expressions obtained via our central Theorem 8. We insist on the fact that such an IRLS algorithm could not be developed with the flagification of pennec_barycentric_2018 and mankovich_fun_2024, since a sum of square roots does not correspond to a least squares problem.
@@
In this work, we explore one specific problem of RSR for conciseness, but we could investigate many other related problems, including Robust PCA. Notably, drawing from the Grassmann averages (GA) method (hauberg_scalable_2016), one could develop many new multilevel RSR and Robust PCA objective functions. The idea behind GA is to replace data points with 1D subspaces ($\mathcal{S}_{i}=\operatorname{Span}(x_{i})$) and then perform subspace averaging methods to find a robust prototype for the dataset. GA ends up solving problems of the form $\operatorname{argmin}_{\mathcal{S}\in\operatorname{Gr}(p,1)}\sum_{i=1}^{n}w_{i}\,\operatorname{dist}_{\operatorname{Gr}(p,1)}^{2}(\operatorname{Span}(x_{i}),\mathcal{S})$, where $w_{i}$ are some weights and $\operatorname{dist}_{\operatorname{Gr}(p,1)}$ is a particular subspace distance detailed in hauberg_scalable_2016. Using instead some multidimensional subspace distances, like the principal angles and its variants (hamm_grassmann_2008; ye_schubert_2016), we can develop many variants of the Grassmann averages, of the form $\operatorname{argmin}_{\mathcal{S}_{1:d}\in\operatorname{Gr}(p,q)}\sum_{i=1}^{n}w_{i}\,\rho(\sqrt{{x_{i}}^{\top}\Pi_{\mathcal{S}}{x_{i}}})$, where $\rho\colon\mathbb{R}\to\mathbb{R}$ is a real function, like $\rho(x)=\arccos(x)$ if we want subspace-angle-like distances, $\rho(x)=-x^{2}$ if we want PCA-like solutions, and many other possible robust variants. Applying the flag trick to those problems yields the following robust multilevel problem: $\operatorname{argmin}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\sum_{i=1}^{n}w_{i}\rho(\sqrt{{x_{i}}^{\top}\Pi_{\mathcal{S}_{1:d}}{x_{i}}})$.
@@ 4.3 The Flag Trick for Trace Ratio Problems
As shown in introduction and particularly in Figure 1, trace ratio methods based on Grassmannian optimization such as linear discriminant analysis might suffer from the nestedness issue. This notably yields inconsistent data representations across different dimensions, as seen on the middle part of the figure, where the clusters get rotated and reflected, leading to cluster identification issues. In this subsection, we address the nestedness issue by applying the flag trick (Definition 10) to trace ratio problems (ngo_trace_2012).
Trace ratio (TR) problems are ubiquitous in machine learning (ngo_trace_2012):
$\operatorname*{arg\,max}_{U\in\operatorname{St}(p,q)}\frac{\operatorname{tr}\left(U^{\top}AU\right)}{\operatorname{tr}\left(U^{\top}BU\right)},$ (16)
@@
where $A,B\in\mathbb{R}^{p\times p}$ are positive semi-definite matrices, with $\operatorname{rank}(B)>p-q$. A famous example of trace ratio problem is Fisher’s linear discriminant analysis (LDA) (fisher_use_1936; belhumeur_eigenfaces_1997). It is common in machine learning to project the data onto a low-dimensional subspace before fitting a classifier, in order to circumvent the curse of dimensionality. It is well known that performing an unsupervised dimension reduction method like PCA comes with the risks of mixing up the classes, since the directions of maximal variance are not necessarily the most discriminating ones (chang_using_1983). The goal of LDA is to use the knowledge of the data labels to learn a linear subspace that does not mix the classes. Let $\penalty 10000\ {X\coloneqq\begin{bmatrix}x_{1}&\cdots&x_{n}\end{bmatrix}\in\mathbb{R}^{p\times n}}$ be a dataset with labels $\penalty 10000\ {Y\coloneqq\begin{bmatrix}y_{1}&\cdots&y_{n}\end{bmatrix}\in{\{1,\dots,C\}}^{n}}$. Let $\mu=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ be the dataset mean and $\mu_{c}=\frac{1}{\#\{i\colon y_{i}=c\}}\sum_{i\colon y_{i}=c}x_{i}$ be the class-wise means. The idea of LDA is to search for a subspace $\mathcal{S}\in\operatorname{Gr}(p,q)$ that simultaneously maximizes the projected between-class variance $\sum_{c=1}^{C}{\|\Pi_{\mathcal{S}}\mu_{c}-\Pi_{\mathcal{S}}\mu\|}_{2}^{2}$ and minimizes the projected within-class variance $\sum_{c=1}^{C}\sum_{i\colon y_{i}=c}{\|\Pi_{\mathcal{S}}x_{i}-\Pi_{\mathcal{S}}\mu_{c}\|}_{2}^{2}$. This can be reformulated as a trace ratio problem (16), with $A=\sum_{c=1}^{C}(\mu_{c}-\mu)(\mu_{c}-\mu)^{\top}$ and $B=\sum_{c=1}^{C}\sum_{i\colon y_{i}=c}(x_{i}-\mu_{c})(x_{i}-\mu_{c})^{\top}$.
@@
More generally, a large family of dimension reduction methods can be reformulated as a TR problem. The seminal work of yan_graph_2007 shows that many dimension reduction and manifold learning objective functions can be written as a trace ratio involving Laplacian matrices of attraction and repulsion graphs. Intuitively, those graphs determine which points should be close in the latent space and which ones should be far apart. Other methods involving a ratio of traces are multi-view learning (wang_trace_2023), partial least squares (PLS) (geladi_partial_1986; barker_partial_2003) and canonical correlation analysis (CCA) (hardoon_canonical_2004), although these methods are originally sequential problems (cf. Remark 1) and not subspace problems.
@@
Classical Newton-like algorithms for solving the TR problem (16) come from the seminal works of guo_generalized_2003; wang_trace_2007; jia_trace_2009. The interest of optimizing a trace-ratio instead of a ratio-trace (of the form $\operatorname{tr}\left((U^{\top}BU)^{-1}(U^{\top}AU)\right)$)—which enjoys an explicit solution given by a generalized eigenvalue decomposition—is also tackled in those papers. The repulsion Laplacians (kokiopoulou_enhanced_2009) instead propose to solve a regularized version $\operatorname{tr}\left(U^{\top}BU\right)-\rho\operatorname{tr}\left(U^{\top}AU\right)$, which enjoys an explicit solution but now has a hyperparameter $\rho$—whereas the latter is directly optimized in the previous approaches.